Five years ago,a father was 3 times as old as his son.Now their combined ages amount to 110 years.find the present age of the father.

Answer by ankor@dixie-net.com(22740)

Five years ago,a father was 3 times as old as his son.
f - 5 = 3(s - 5)
f - 5 = 3s - 15
f = 3s - 15 + 5
f = 3s - 10
Now their combined ages amount to 110 years.
f + s = 110
Replace f with (3s-10)
(3s-10 + s = 110
3s + s = 110 + 10
4s = 120
s = 120/4
s = 30 yrs is the son
:
find the present age of the father. I'll let you do this